OpenStudy (anonymous):

Solve Wc = Tp - 0.8Wf Where Wc = 0 and Tp = 16

OpenStudy (anonymous):

substitute the numbers in the equation

OpenStudy (anonymous):

I did, I'm stuck

OpenStudy (anonymous):

Wc = Tp - 0.8 * Wf 0 = 16 - 0.8 * Wf stuck at this step?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

0 = 16 - 0.8Wf +0.8Wf both sides 0.8Wf = 16 divide by 0.8 both sides Wf = 20

OpenStudy (anonymous):

Oh I see thank you!